OpenStudy (anonymous):

Which statements are true about the value of the digits in the number 4,444,444? Choose all answers that are correct. A. The digit in the hundred thousands place is 10 times the value of the digit in the ten thousands place. B. The digit in the thousands place is 40 times the value of the digit in the hundreds place. C. The digit in the thousands place is 1/4 the value of the digit in the ten thousands place. D. The digit in the ten thousands place is 1/10 the value of the digit in the hundred thousands place.

OpenStudy (anonymous):

help me will give u a medal

OpenStudy (texaschic101):

A. digit in the 10,000 place is : 40,000 digit in the 100,000 place is : 400,000 40,000 * 10 = 400,000 (This one is TRUE) B. digit in 1000 place is : 4000 digit in the 100 place is : 400 400 * 40 = 16,000 (This one would be FALSE) C. digit in 1000 place is : 4000 digit in 10,000 place is : 40,000 1/4 of 40,000 = 40,000/4 = 10,000 (This one is FALSE) D. digit in 10,000 place is : 40,000 digit in 100,000 place is : 400,000 1/10 of 400,000 = 400,000/10 = 40,000 (This one is TRUE)